# Functional-level Uncertainty Quantification for Calibrated Fine-tuning on LLMs

### v3

@@ 1 Introduction
During inference, LoRA experts offer diverse functional relationships acquired during fine-tuning, while MoE routers dynamically select the most relevant experts for each input prompt. This selection is guided by functional-level uncertainty calibration performed throughout fine-tuning, which aims to optimize the choice of the correct functional relationship for each prompt. More accurate expert selection enables the model to learn diverse functional relationships. As a result, the model’s uncertainty estimates become better aligned, enhancing calibration without compromising accuracy. To summarize, our contributions include:
@@ 2 Preliminaries
Here, $\Delta\mathbf{W}$ is calculated as the product of two matrices, $\mathbf{B}\in\mathbb{R}^{n_{\text{out}}\times n_{\text{lr}}}$ and $\mathbf{A}\in\mathbb{R}^{n_{\text{lr}}\times n_{\text{in}}}$ where $n_{\text{lr}}$ is significantly smaller than $n_{\text{in}}$ or $n_{\text{out}}$. For example, we use $n_{\text{lr}}=32$ while $n_{\text{in}}=n_{\text{out}}=4096$ for the Llama2-7b model (Touvron et al., 2023a). Therefore, the total number of LoRA parameters for this $\Delta\mathbf{W}$ is $n_{\text{lr}}(n_{\text{in}}+n_{\text{out}})$, which is far smaller than the parameter count of the full matrix, $n_{\text{in}}n_{\text{out}}$. One of the key motivations of incorporating LoRA to fine-tune LLMs is the vast amount of memory cost reduction compared with fine-tuning on the full model. For an LLM with $7$ billion parameters, maintaining the average gradient and average squared gradients for optimization multiplies the memory required by a factor of $3$ compared to simply loading model weights. LoRA greatly mitigates this memory cost as the tripled memory consumption only applies to LoRA adapters.
@@
Let $\mathbf{h}^{\ell}\in\mathbb{R}^{1\times d}$ $(1\leq\ell\leq L)$ denote the output hidden state at the $\ell$-th layer of the LLM, where $L$ is the number of LLM layers and $d$ is the hidden dimension. With $\mathbf{W}_{r}^{\ell}$ as the trainable router weight at layer $\ell$, the top-k gate router $\tilde{R}(\cdot)$ chooses $k$ experts with highest probability given a hidden state $\mathbf{h}^{\ell}$:
@@ 3.2 Quantifying FLU with MixLoRA
In practice, as illustrated in Figure 2, the top-$k$ router at each layer produces a sparse probability vector $\alpha^{\ell}=(\alpha_{1}^{\ell},...,\alpha_{k}^{\ell})$, dynamically mixing the basis functions captured by the LoRA experts given the current hidden state. The values of the top-$K$ routing weights that contribute to the final output hidden state serve as a direct quantification of the model’s uncertainty at functional-level. We follow the routing mechanisms used in MoE layers (see Eq. (3) and (4)), employing top-$2$ gate routers for the mixture. At each layer $\ell$, we compute the raw router probabilities and retrain the largest two:
@@ 3.3 Calibration Loss
The first term is an indicator function that equals $1$ if the model prediction matches the ground truth $y^{*}$ for prompt $x$, and 00 otherwise, corresponding to a one-hot definition of ground truth confidence.
This calibration loss directly encourages the functional-level uncertainty (FLU) to reflect the true correctness of the model’s predictions. As shown in Figure 2, when the mixture model predicts correctly, the loss pushes FLU toward $1$ (high confidence); when incorrect, toward 00 (low confidence). In other words, when optimized over the data distribution, the calibration loss pushes FLU to represent the probability of prediction correctness. We formally state this property as follows and present the proof in Appendix A.1:
@@
¯$\bar{\mathcal{L}}_{\mathrm{cal}}=\mathbb{E}_{(x,y^{*})\sim\mathcal{D}}\left(\mathbbm{1}\{\mathrm{MixLoRA}(x)=y^{*}\}-\mathrm{FLU}(x)\right)^{2}.$
@@ 3.4 Fine-tuning with the Total Loss
$\mathcal{L}=\mathbf{CE}+\gamma\cdot\mathcal{L}b+\beta\cdot\mathcal{L}_{\mathrm{cal}},$ (14)
where $\mathbf{CE}$ is the cross-entropy loss, and $\gamma$, $\beta$ are hyperparameters for the auxiliary terms. Details of $\mathcal{L}_{b}$ are provided in Appendix A.2.
@@ 4 Related Work
Large Language Models (LLMs) have achieved impressive performance across diverse NLP tasks (Brown et al., 2020; Hoffmann et al., 2022; Touvron et al., 2023b, c), with instruction fine-tuning (Chung et al., 2022; Iyer et al., 2022; Zheng et al., 2024) further boosting their adaptability for conversational AI (Wu et al., 2023b; Achiam et al., 2023). However, scaling LLMs increases the resource demands of full fine-tuning. Parameter-efficient fine-tuning (PEFT) methods (Mangrulkar et al., 2022)—such as LoRA (Hu et al., 2021b) and its variants (Kopiczko et al., 2023; Hyeon-Woo et al., 2021; Renduchintala et al., 2023; Zhang et al., 2023; Liu et al., 2024)—reduce adaptation costs by updating a subset of parameters.
@@
Established uncertainty quantification methods have been studied in conjunction with the LoRA structure for LLMs. Monte-Carlo dropout (Gal and Ghahramani, 2016) interprets dropout in neural networks as approximate Bayesian inference in deep Gaussian processes, allowing uncertainty estimates to be obtained from existing LoRA adapters without modifying them. Checkpoint ensemble (Chen et al., 2017) utilizes predictions from multiple LoRA checkpoints saved during a single fine-tuning process to calibrate uncertainty. Deep ensemble (Lakshminarayanan et al., 2017; Wang et al., 2023; Zhai et al., 2023) combines the predictions from multiple LoRA adapters for better uncertainty calibration. Laplace-LoRA (Yang et al., 2024a) applies Bayesian inference via Laplace approximation to the LoRA parameters after fine-tuning, resulting in improved calibration and uncertainty estimates.
@@ 5 Experiments
Figure 3: Performance comparison of different methods fine-tuned with LlaMA2-7B across four common sense reasoning tasks and a domain-specific task. UQ4CT shows substantial ECE improvements while maintaining high accuracy.
@@
We implement UQ4CT with PyTorch (Paszke et al., 2019), extending the MixLoRA repository in (Li et al., 2024). We use the LLaMA-2-7B-hf (Touvron et al., 2023a) as our base model. In particular, we apply MixLoRA to query, key, value and output layers, together with the feed-forward networks in LLaMA-2-7B-hf (gate layer, down layer and up layer). Details are provided in Appendix A.6.
  • LoRA (Hu et al., 2021a). We use standard LoRA fine-tuning as lower performance bound.
  • Monte Carlo (MC) Dropout (Gal and Ghahramani, 2016) keeps dropout on at both training and testing time. By performing multiple forward passes, MC dropout randomly shuts down a portion of model nodes, producing ensemble-alike predictions. To combine LoRA fine-tuning with MC dropout, we apply dropout on the input of the LoRA adapter, following the implementation of Mangrulkar et al. (2022).
  • Deep Ensemble (Lakshminarayanan et al., 2017) averages the predictions from each ensemble member which have been trained with varying random initialization. We combine deep ensemble with LoRA by fine-tuning $3$ randomly initialized LoRAs together and ensembling their output as final predictions.
  • Laplace-LoRA (LA) (Yang et al., 2024a) applies a post-hoc Laplace approximation on fine-tuned LoRA parameters for better uncertainty estimation.
  • MixLoRA (Li et al., 2024) incorporates LoRAs via a router network to reduce computational overhead while expanding model capacity. We add this as a baseline to resemble plain LoRA MoE model performance.
@@
To assess robustness under distribution shifts, we fine-tune models on the OBQA dataset and evaluate them following (Yang et al., 2024a). We use ARC-C and ARC-E to represent smaller distribution shifts, as these datasets focus on general science reasoning similar to OBQA but are more challenging and diverse. For larger shifts, we utilize the four aforementioned domain-specific MMLU subtasks, which span a wide range of expertise from elementary to professional levels. This domain specificity represents a greater distribution shift from the general common sense focus of OBQA.
@@
Table 1: Performance comparison of different methods fine-tuned on the OBQA dataset with LlaMA2-7B across two smaller distribution shift (DS) tasks and four larger distribution shift tasks. UQ4CT shows substantial ECE improvements while maintaining high accuracy.
@@
ACC $\uparrow$ LoRA $72.1_{0.87}$ $58.6_{1.93}$ $66.5_{3.38}$ $35.5_{2.35}$ $30.8_{1.72}$ $34.9_{1.41}$ $39.1_{1.52}$
MC Drop $74.8_{1.34}$ $58.7_{2.07}$ $66.6_{3.30}$ $36.0_{1.69}$ $30.3_{2.25}$ $35.1_{0.86}$ $39.1_{1.35}$
Ensemble $75.5_{1.4}$ $57.7_{0.78}$ $69.1_{0.48}$ $\mathbf{36.7_{2.18}}$ $30.3_{1.13}$ $35.3_{1.02}$ $39.9_{1.89}$
LA $70.8_{1.24}$ $58.7_{0.58}$ $67.9_{0.41}$ $33.7_{1.22}$ $29.6_{1.32}$ $35.4_{0.75}$ $38.5_{1.61}$
MixLoRA $75.5_{2.91}$ $58.5_{1.44}$ $\mathbf{69.2_{1.02}}$ $35.2_{2.92}$ $30.3_{0.98}$ $\mathbf{35.9_{0.43}}$ $\mathbf{40.6_{1.13}}$
UQ4CT $\mathbf{77.3_{1.36}}$ $\mathbf{58.8_{1.06}}$ $65.8_{1.31}$ $\underline{36.2_{1.24}}$ $\mathbf{34.1_{2.31}}$ $\underline{35.8_{1.01}}$ $\underline{40.0_{1.24}}$
ECE $\downarrow$ LoRA $10.2_{1.07}$ $16.7_{2.28}$ $13.3_{2.48}$ $29.7_{2.69}$ $32.3_{1.85}$ $29.2_{3.08}$ $31.0_{2.13}$
MC Drop $10.9_{0.24}$ $16.7_{2.20}$ $13.2_{2.21}$ $23.2_{2.32}$ $31.6_{1.64}$ $28.0_{2.93}$ $25.9_{2.27}$
Ensemble $8.83_{2.35}$ $15.1_{1.09}$ $11.1_{0.99}$ $22.4_{1.32}$ $28.5_{2.13}$ $29.0_{1.37}$ $24.5_{0.39}$
LA $17.2_{1.2}$ $16.2_{0.5}$ $24.4_{0.42}$ $28.6_{2.61}$ $30.5_{1.43}$ $29.5_{1.83}$ $30.7_{1.2}$
MixLoRA $12.9_{1.99}$ $19.0_{1.88}$ $14.5_{2.57}$ $26.4_{3.25}$ $33.7_{1.87}$ $30.3_{2.27}$ $28.3_{1.07}$
UQ4CT $\mathbf{5.0_{1.15}}$ $\mathbf{8.9_{3.46}}$ $\mathbf{6.5_{1.85}}$ $\mathbf{19.6_{2.90}}$ $\mathbf{23.1_{1.17}}$ $\mathbf{25.9_{3.43}}$ $\mathbf{21.9_{3.49}}$

### v4

@@ 1 Introduction
During inference, LoRA experts offer diverse functional relationships acquired during fine-tuning, while MoE routers dynamically select the most relevant experts for each input prompt. This selection is guided by functional-level uncertainty calibration performed throughout fine-tuning, which aims to optimize the choice of the correct functional relationship for each prompt. More accurate expert selection enables the model to learn diverse functional relationships. As a result, the model’s uncertainty estimates become better aligned, enhancing calibration without compromising accuracy.
To summarize, our contributions include:
@@ 2 Preliminaries
Here, $\Delta\mathbf{W}$ is calculated as the product of two matrices, $\mathbf{B}\in\mathbb{R}^{n_{\text{out}}\times n_{\text{lr}}}$ and $\mathbf{A}\in\mathbb{R}^{n_{\text{lr}}\times n_{\text{in}}}$ where $n_{\text{lr}}$ is significantly smaller than $n_{\text{in}}$ or $n_{\text{out}}$. For example, we use $n_{\text{lr}}=32$ while $n_{\text{in}}=n_{\text{out}}=4096$ for the Llama3.1-8B model (Dubey et al., 2024). Therefore, the total number of LoRA parameters for this $\Delta\mathbf{W}$ is $n_{\text{lr}}(n_{\text{in}}+n_{\text{out}})$, which is far smaller than the parameter count of the full matrix, $n_{\text{in}}n_{\text{out}}$. One of the key motivations of incorporating LoRA to fine-tune LLMs is the vast amount of memory cost reduction compared with fine-tuning on the full model. For an LLM with $7$ billion parameters, maintaining the average gradient and average squared gradients for optimization multiplies the memory required by a factor of $3$ compared to simply loading model weights. LoRA greatly mitigates this memory cost as the tripled memory consumption only applies to LoRA adapters.
@@
[…]
@@ 3.2 Quantifying FLU with MixLoRA
In practice, as illustrated in Figure 2, the top-$k$ router at each layer produces a sparse probability vector $\alpha^{\ell}=(\alpha_{1}^{\ell},...,\alpha_{k}^{\ell})$, dynamically mixing the basis functions captured by the LoRA experts given the current hidden state. The values of the top-$K$ routing weights that contribute to the final output hidden state serve as a direct quantification of the model’s uncertainty at functional-level. We follow the routing mechanisms used in MoE layers (see Eq. (3) and (4)), employing top-$2$ gate routers for the mixture. At each layer $\ell$, we compute the raw router probabilities and retain the largest two:
@@ 3.3 Calibration Loss
The first term is an indicator function that equals $1$ if the model prediction matches the ground truth $y^{*}$ for prompt $x$, and 0 otherwise, corresponding to a one-hot definition of ground truth confidence.
This calibration loss directly encourages the functional-level uncertainty (FLU) to reflect the true correctness of the model’s predictions. As shown in Figure 2, when the mixture model predicts correctly, the loss pushes FLU toward $1$ (high confidence); when incorrect, toward 0 (low confidence). In other words, when optimized over the data distribution, the calibration loss pushes FLU to represent the probability of prediction correctness. We formally state this property as follows and present the proof in Appendix A.1:
@@
$\bar{\mathcal{L}}_{\mathrm{cal}}=\mathbb{E}_{(x,y^{*})\sim\mathcal{D}}\left(\mathbbm{1}\{\mathrm{MixLoRA}(x)=y^{*}\}-\mathrm{FLU}(x)\right)^{2}.$
@@ 3.4 Fine-tuning with the Total Loss
$\mathcal{L}=\mathbf{CE}+\gamma\cdot\mathcal{L}_{b}+\beta\cdot\mathcal{L}_{\mathrm{cal}},$ (14)
where $\mathbf{CE}$ is the cross-entropy loss, and $\gamma$, $\beta$ are hyperparameters for the auxiliary terms. We fix $\gamma$, $\beta$ to $1$ for our experiments. Details of $\mathcal{L}_{b}$ are provided in Appendix A.2.
@@ 4 Related Work
Large Language Models (LLMs) have achieved impressive performance across diverse NLP tasks (Brown et al., 2020; Hoffmann et al., 2022; Touvron et al., 2023a, b), with instruction fine-tuning (Chung et al., 2022; Iyer et al., 2022; Zheng et al., 2024) further boosting their adaptability for conversational AI (Wu et al., 2023b; Achiam et al., 2023). However, scaling LLMs increases the resource demands of full fine-tuning. Parameter-efficient fine-tuning (PEFT) methods (Mangrulkar et al., 2022)—such as LoRA (Hu et al., 2021b) and its variants (Kopiczko et al., 2023; Hyeon-Woo et al., 2021; Renduchintala et al., 2023; Zhang et al., 2023; Liu et al., 2024)—reduce adaptation costs by updating a subset of parameters.
@@
Established uncertainty quantification methods have been studied in conjunction with the LoRA structure for LLMs. Monte-Carlo dropout (Gal and Ghahramani, 2016) interprets dropout in neural networks as approximate Bayesian inference in deep Gaussian processes, allowing uncertainty estimates to be obtained from existing LoRA adapters without modifying them. Checkpoint ensemble (Chen et al., 2017) utilizes predictions from multiple LoRA checkpoints saved during a single fine-tuning process to calibrate uncertainty. Deep ensemble (Lakshminarayanan et al., 2017; Wang et al., 2023; Zhai et al., 2023) combines the predictions from multiple LoRA adapters for better uncertainty calibration. Laplace-LoRA (Yang et al., 2024a) applies Bayesian inference via Laplace approximation to the LoRA parameters after fine-tuning, resulting in improved calibration and uncertainty estimates. Bayesian Low-Rank Adaptation by Backpropagation (BLoB) (Wang et al., 2024) extends the LA method by jointly optimizing the mean and covariance of LoRA parameters via backpropagation throughout fine-tuning.
@@ 5 Experiments
Table 1: Performance comparison of different methods fine-tuned with LlaMA3.1-8B across four common sense reasoning tasks and a domain-specific task. UQ4CT shows substantial ECE improvements while maintaining high accuracy.
Metrics Methods BoolQ ARC-E ARC-C OBQA ClimateQA
ACC $\uparrow$ LoRA $\underline{89.73_{0.58}}$ $\underline{88.82_{1.82}}$ $78.21_{0.73}$ $88.00_{1.22}$ $78.25_{1.29}$
MC Drop $9.65_{0.55} 88$$8.16_{1.75}$ $77.14_{0.69}$ $87.12_{1.18}$ $78.19_{1.35}$
Ensemble $\mathbf{89.87_{0.43}}$ $\mathbf{89.14_{1.31}}$ $78.81_{0.96}$ $86.47_{0.42}$ $78.53_{2.98}$
MixLoRA $88.68_{0.82}$ $87.74_{0.36}$ $78.56_{1.87}$ $\underline{88.27_{0.50}}$ $\underline{79.94_{1.29}}$
LA $89.58_{0.19}$ $86.22_{3.52}$ $78.00_{3.76}$ $86.00_{6.01}$ $79.82_{3.48}$
BLoB(Mean) $9.02_{0.93} 88$$8.71_{0.82}$ $79.37_{0.71}$ $87.60_{1.04}$ $79.02_{0.50}$
BLoB(N=10) $89.39_{1.13}$ $87.96_{0.62}$ $\mathbf{80.08_{1.55}}$ $87.13_{0.88}$ $79.02_{0.50}$
UQ4CT $9.17_{1.33} 88$$8.66_{0.20}$ $\underline{79.60_{1.21}}$ $\mathbf{88.40_{0.35}}$ $\mathbf{79.97_{0.85}}$
ECE $\downarrow$ LoRA $15.82_{0.57}$ $6.55_{1.70}$ $14.07_{0.68}$ $7.30_{0.43}$ $13.70_{1.50}$
MC Drop $14.73_{0.54}$ $6.48_{1.74}$ $14.12_{0.71}$ $7.24_{0.39}$ $13.11_{1.46}$
Ensemble $14.56_{0.55}$ $7.08_{0.73}$ $13.71_{1.29}$ $8.63_{0.38}$ $14.69_{0.84}$
MixLoRA $15.85_{0.76}$ $7.79_{0.45}$ $13.71_{1.90}$ $6.58_{0.21}$ $14.68_{0.09}$
LA $3.78_{0.60}$ $7.63_{1.71}$ $8.92_{4.16}$ $1.97_{5.97} 11$$1.48_{1.66}$
BLoB(Mean) $7.54_{0.57}$ $4.89_{0.32}$ $11.26_{1.13}$ $6.83_{0.90}$ $12.74_{0.88}$
BLoB(N=10) $\underline{2.76_{0.41}}$ $\mathbf{3.35_{0.50}}$ $\underline{6.81_{1.43}}$ $\underline{3.84_{1.00}}$ $\underline{11.96_{2.57}}$
UQ4CT $\mathbf{1.79_{0.43}}$ $\underline{3.97_{0.78}}$ $\mathbf{4.43_{0.82}}$ $\mathbf{3.34_{1.60}}$ $\mathbf{9.36_{2.77}}$
@@
We implement UQ4CT with PyTorch (Paszke et al., 2019), extending the MixLoRA repository in (Li et al., 2024). We use the Llama-3.1-8B (Touvron et al., 2023c) as our base model. In particular, we apply MixLoRA to query, key, value and output layers, together with the feed-forward networks in LLaMA-3.1-8B (gate layer, down layer and up layer). Details are provided in Appendix A.6.
Baselines.
We compare UQ4CT with state-of-the-art uncertainty estimation methods along with naive fine-tuning applied to the LoRA adapters of LLMs, including LoRA (Hu et al., 2021a), Monte Carlo (MC) Dropout (Gal and Ghahramani, 2016), Deep Ensemble (Lakshminarayanan et al., 2017), Laplace-LoRA (LA) (Yang et al., 2024a), Bayesian Low-Rank Adaptation by Backpropagation (BLoB) (Wang et al., 2024) and MixLoRA (Li et al., 2024). Note that for BLoB(N=10), the method performs $10$ forward passes with differently sampled LoRA parameters for each question, which is a unfair computational budget advantage compared against UQ4CT with only $1$ forward pass.
@@
To assess robustness under distribution shifts, we fine-tune models on the OBQA dataset and evaluate them following Yang et al. (2024a). We use ARC-C and ARC-E to represent smaller distribution shifts, as these datasets focus on general science reasoning similar to OBQA but are more challenging and diverse. For larger shifts, we utilize the four aforementioned domain-specific MMLU subtasks, which span a wide range of expertise from elementary to professional levels. This domain specificity represents a greater distribution shift from the general common sense focus of OBQA.
@@
Table 2: Performance comparison of different methods fine-tuned on the OBQA dataset with LlaMA3.1-8B across two smaller distribution shift (DS) tasks and four larger distribution shift tasks. UQ4CT shows substantial ECE improvements while maintaining high accuracy.
@@
ACC $\uparrow$ LoRA $88.0_{1.22}$ $77.8_{0.16}$ $86.7_{0.77}$ $\mathbf{55.8_{0.52}}$ $54.3_{3.30}$ $44.9_{0.23}$ $58.8_{0.21}$
MC Drop $87.1_{1.18}$ $77.1_{2.01}$ $86.9_{2.42}$ $54.4_{1.58}$ $54.1_{1.82}$ $45.0_{0.76}$ $58.3_{1.46}$
Ensemble $86.5_{0.42}$ $78.2_{0.90}$ $85.4_{0.47}$ $53.8_{1.02}$ $52.4_{0.56}$ $45.0_{0.20}$ $60.6_{0.57}$
MixLoRA $\underline{88.3_{0.50}}$ $78.1_{0.45}$ $86.7_{0.35}$ $53.1_{1.14}$ $\underline{54.7_{2.28}}$ $45.0_{1.46}$ $\underline{60.9_{1.04}}$
LA $86.0_{6.01}$ $78.7_{0.55}$ $86.4_{0.76}$ $\underline{54.7_{1.82}}$ $53.6_{2.77}$ $44.9_{1.03}$ $59.7_{0.94}$
BLoB(Mean) $87.6_{1.04}$ $\underline{79.5_{1.10}}$ $86.6_{0.65}$ $51.2_{0.99}$ $48.6_{1.44}$ $39.9_{7.85}$ $57.0_{3.51}$
BLoB(N=10) $87.1_{0.88}$ $\mathbf{79.8_{1.06}}$ $\underline{87.2_{0.79}}$ $52.8_{1.28}$ $51.9_{3.13}$ $43.8_{4.60}$ $58.5_{5.33}$
UQ4CT $\mathbf{88.4_{0.35}}$ $79.0_{0.56}$ $\mathbf{87.8_{0.47}}$ $53.3_{0.61}$ $\mathbf{61.1_{3.20}}$ $\mathbf{45.4_{0.50}}$ $\mathbf{61.1_{1.48}}$
ECE $\downarrow$ LoRA $7.30_{0.43}$ $14.8_{0.62}$ $9.6_{0.69}$ $21.0_{3.05}$ $24.1_{3.63}$ $29.3_{1.98}$ $24.0_{1.81}$
MC Drop $7.24_{0.39}$ $13.4_{2.15}$ $10.2_{1.89}$ $20.8_{3.26}$ $24.1_{0.77}$ $29.1_{0.64}$ $21.6_{3.92}$
Ensemble $8.63_{0.38}$ $15.4_{0.46}$ $10.7_{0.56}$ $14.0_{3.18}$ $17.4_{1.98}$ $19.9_{2.95}$ $16.1_{2.07}$
MixLoRA $6.58_{0.21}$ $14.5_{0.55}$ $9.9_{0.20}$ $17.1_{3.06}$ $17.8_{2.80}$ $21.6_{4.07}$ $18.0_{2.78}$
LA $11.97_{5.97}$ $7.2_{0.5}$ $6.4_{0.42}$ $13.7_{2.14}$ $15.5_{2.0}$ $\underline{19.0_{0.54}}$ $15.7_{2.30}$
BLoB(Mean) $6.83_{0.90}$ $11.37_{1.94}$ $6.6_{1.65}$ $17.2_{2.72}$ $18.5_{2.82}$ $22.6_{1.26}$ $16.9_{3.02}$
BLoB(N=10) $\underline{3.84_{1.00}}$ $\underline{5.8_{0.96}}$ $\mathbf{3.0_{0.87}}$ $\underline{11.5_{2.76}}$ $\underline{14.9_{1.93}}$ $19.7_{3.21}$ $\underline{14.5_{3.38}}$
UQ4CT $\mathbf{3.34_{1.60}}$ $\mathbf{3.6_{1.44}}$ $\underline{3.6_{1.32}}$ $\mathbf{10.8_{3.73}}$ $\mathbf{13.2_{1.86}}$ $\mathbf{18.1_{4.40}}$ $\mathbf{13.2_{4.06}}$